20. Februar 2018: Dienstag

Today’s Activities:

  • Backups:
    • Reclaimed six (6) tapes for the vaults over the long weekend.
    • Reclaimed four (4) tapes for the vaults.
    • Replaced bujbod2’s 8:21.
      I had done the pdprprmv yesterday which completed.
      Therefore (SURPRISE) it did COPYBACK to the previous HSP!
    • Clones started for the week.
    • COMIS clones complete immediately, having found nothing to do.
    • Calculate daily change rates of some of the systems (changeRate directory).
    • TAPE Clones completed at 17:11.
  • LDAP/Accounts:
    • peregrine (one of the replicas) broke at 11pm Monday night.
      The ESX system it was running on died with the PSOD.
      On Call person did not know how to start up the LDAP service
      after a reboot.
      Updated the wiki entry to specifically spell out the process.
    • There are no issues requiring human intervention this morning.
    • Looks like Banner feed will be changing (getting larger).
  • Footprints:
    • BMC finds problem – replaces quickcreatefile.storable…

Tags:

Leave a Reply

You must be logged in to post a comment.